What is the pero probability that ramdomly selected year will have precipitation greater than 18 inches for the first 7 months?

1
Expert's answer
2021-11-17T03:47:59-0500

Let "X" be a random variable representing the precipitation for the first 7 months of a year. "X" follows a normal distribution with mean "\mu=19.32" and standard deviation "\sigma=2.44". We need to find the probability,

"p(X\gt 18)=p((X-\mu)/\sigma\gt (18-\mu)/\sigma)=p(Z\gt (18-19.32)/2.44)=p(Z\gt -0.5409836)"

This probability can be written as,

"p(Z\gt -0.5409836)=1-p(Z\lt -0.5409836)=1-0.2946=0.7054"

Therefore, the  probability that a randomly selected year will have precipitation greater than 18 inches for the first 7 months is 0.7054.